# Revealing the nature of ultra-long period objects with space-based gravitational-wave interferometers

### v2

@@ 3 Orbital gravitational waves
Assuming that the pulse period is that of the orbit, we use the data collated in Table 1 to estimate $h_{0}$ from equation (7) for each ULP using the chirp mass spread considered in Sec. 2. Notably, the period derivative is moderate in most systems ($|\dot{P}|_{\rm max}\lesssim 10^{-9}\text{ ss}^{-1}$; Coti Zelati & Borghese, 2024), meaning that GW signals should be essentially monochromatic over multi-year observational windows, $T$. This allows for an accumulation of signal power over many orbital cycles, $N=Tf_{\rm GW}=2T/P_{\rm orb}$. Averaging over orbital orientations and polarisations, the characteristic strain $h_{c}$ felt by a detector can thus be estimated through $h_{c}\approx\sqrt{2N}h_{0}$ (Finn & Thorne, 2000), with value
$h_{c}\approx$   $\,\,2.1\times 10^{-20}\left(\frac{1\text{ hr}}{P_{\text{orb}}}\right)^{7/6}\left(\frac{\mathcal{M}}{0.5M_{\odot}}\right)^{5/3}$ (8)
×$\qquad\times\left(\frac{T}{0.5\text{ yr}}\right)^{1/2}\left(\frac{1\text{ kpc}}{d}\right).$
@@
Using the LISA design specifications from Amaro-Seoane et al. (2017), we adopt the noise power spectral density ($S_{n}$) fits from Robson, Cornish & Liu (2019), which include intrinsic noise (from single-link optical metrology and test mass acceleration) as well as Galactic confusion noise from unresolved sources, to estimate ULP detectability. Figure 1 shows characteristic strains $h_{c}$ from equation (8), as functions of GW frequency, for objects listed in Tab. 1 assuming an observation time of two years. We find that two of the known ULPs (solid symbols) could be detectable by LISA (and also by Taiji and TianQin, though not shown), with a couple others being marginal, under favourable conditions (see Sec. 3.1). As it happens, the most promising candidates are also among the most intriguing.
@@
CHIME J0630. Given its proximity to Earth, insights on local supernovae rates could be gleaned if this source is a neutron star, a position which would be firm if no detection is made shortly after LISA launch. The period would thus correspond to rotation, requiring that some magnetars spindown faster than stipulated by pure dipole braking (e.g. via mass-loaded winds; Beniamini, Wadiasingh & Metzger, 2020). An excess of nearby neutron stars (and magnetars in particular) could also call for revisions to population synthesis models (though see Popov et al., 2003). While a nearby white dwarf is not particularly unusual, Tauris (2018) estimates that for SNRs $\gtrsim 100$, $\mathcal{M}$ could be measured to within $\lesssim 1\%$ from which constraints on binary makeup, and possibly even pre-Newtonian effects predicted by some modified theories of gravity (Littenberg & Yunes, 2019), follow.
@@
GX J1627. Discovered in 2018 archival data from the Murchison Widefield Array, this object has since been silent (Hurley-Walker et al., 2022). If interpreted as a neutron star, because the spindown luminosity, $L_{\rm sd}\lesssim 10^{28}$ erg/s, is much lower than the radio luminosity, $L^{1.4\text{GHz}}_{\nu}\sim 4\times 10^{31}$ erg/s, it could not be rotation powered. If magnetically powered, the deep X-ray limits set by Chandra ($L^{\rm max}_{\rm X}\sim 10^{30}$ erg/s; Rea et al., 2022) mean that a non-detection by LISA could indicate that (some) ULPs populate a magnetar branch that is distinct from soft gamma repeaters and anomalous X-ray pulsars (Suvorov, Dehman & Pons, 2025). For a dwarf, the lack of radio reactivation points towards either a choked magnetosphere (e.g. via accretion) or that emissions are beamed inconsistently, either of which holds promise to inform on tight-binary interaction.
@@
DA J1832. Because this ULP has shown contemporaneous X-ray bursting with radio pulsing (Wang et al., 2025), a magnetar progenitor is favoured over a binary (Suvorov, Dehman & Pons, 2025). If confirmed as a binary however, sporadic accretion could be responsible for igniting a burst of comparable luminosity to that seen from DA J1832 in February 2024 ($L_{\rm X}\sim 10^{33}$ erg/s), though it would be difficult to explain the lack of any quiescent emissions (see Mukai, 2017). If the object enters into an outburst phase in future, searches for absorption lines could be useful to constrain its nature. The fact that the source was radio-loud for a $\sim$ year is also uncharacteristic for transitional pulsars (for instance), and so a detection could insist on revisions to theoretical models from a variety of directions.
@@
A J1935. This source displays a unique property amongst ULPs, being that of state switching: there are pulse states with high-degrees of linear polarisation, others which are predominantly circular, and a null state (Caleb et al., 2024). A high degree of linear polarisation and frequent nulling is rather reminiscent of radio-loud magnetars. For an isolated object, a highly-dynamic magnetosphere would be required to explain this temperamental behaviour; the evolution could be driven, for example, by helicity injections from the plastic motions of crustal platelets (e.g. Beloborodov, 2009).
@@ 3.1 Signal-to-noise ratios
In general, a confident detection requires an SNR of at least $\sim$ 10. While detection odds improve for sources with known sky locations, we provide conservative estimates by avoiding complications related to barycentric and other corrections. Borrowing formula (26) from Robson, Cornish & Liu (2019), the orientation- and polarisation-averaged SNR is estimable through
$\text{SNR}\approx\frac{8\pi^{2/3}G^{5/3}\mathcal{M}^{5/3}f_{\rm GW}^{2/3}}{\sqrt{5}c^{4}d}\sqrt{\frac{T}{S_{n}(f_{\rm GW})}}\ ,$ (9)
for speed of light $c$ and Newton’s constant $G$. We calculate expression (9) for sources listed in Tab. 1. The results are shown in Table 2 for two scenarios: (i) using the maximum DM-inferred distances and lowest values of $\mathcal{M}$ considered in Sec. 2, and (ii) an optimistic case where the masses and distances take their upper and lower confidence bounds, respectively. The latter scenario is indicated by bracketed values in the second ($T=0.5$ yr) and third ($T=4$ yr) columns of Tab. 2.
As expected visually from Fig. 1, taking $T\lesssim 0.5$ yr practically guarantees that CHIME J0630 and GX J1627 would be identifiable in the LISA data stream if $P=P_{\rm orb}$. While hopes for other ULPs are low, there is some call for optimism for DA J1832 and A J1935 if the systems contain heavier stars, allowing them to reach significant SNRs when accounting for their sky locations. For instance, if A J1935 contains a neutron star and a reasonably heavy companion, its well-constrained sky location could allow for a boost to the SNR by a factor $\sim 2$, which would give a final value of $\text{SNR}(4\text{yr})\sim 12$ which would be sufficient to claim a confident detection. Combining data from LISA with Taiji and TianQin could also lead to non-negligible boosts in SNR for this and other sources, perhaps enabling detections in marginal scenarios like a WDWD for DA J1832.
Table 2: SNRs for LISA for the systems listed in Tab. 1 for $T=0.5$ yr (middle column) or $T=4$ yr (right) in either the least (largest distances and smallest chirp masses) or most (bracketed values) optimistic scenarios.
Source name SNR (0.5 yr; max) SNR (4 yr; max)
@@
ASKAP J1839 $<10^{-3}$ (0.004) $<10^{-3}$ (0.01)
@@ 4 Conclusions
ULPs remain mysterious, with some studies favouring isolated neutron stars and others favouring WDMDs or compact binaries (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). We have argued here that if the pulsation period in ULPs matches that of the orbit or close to it, at least a few should be visible to LISA and its sister spacecraft (Tab. 2). If electromagnetic data remain inconclusive, GW observations could help unveil their nature. Importantly, among the known ULPs, those that may be brightest in the LISA band correspond to objects with uncertain classification (Fig. 1).
While some implications of (non-)detections were highlighted in Secs. 2 and 3, a wealth of physical processes may operate in ULPs. For instance, linking radio luminosities to electromotive losses implies a minimum strength for the primary’s magnetic field. Assuming a maser-like emission mechanism, some systems appear to require GigaGauss strengths (Qu & Zhang, 2025). While dynamo action may amplify an internal field (Schreiber et al., 2021), the magnetospheric field may not reflect the interior strength because magnetic burial – a process where field lines are equatorially advected by infalling plasma – will reduce the global dipole moment during accretion while creating strong multipoles (Suvorov & Melatos, 2020). Typical accretion rates in tight-binary polars are $\sim 10^{-10}M_{\odot}\text{ yr}^{-1}$ (Pala et al., 2020) implying that, for a C/O dwarf with $M_{\star}\sim 0.8M_{\odot}$, the accretion timescale matches that of Ohmic diffusion ($\tau_{\Omega}$) at depths corresponding to densities of $\sim 10^{5}\text{ g cm}^{-3}$ where Cumming (2002) estimates $\tau_{\Omega}\lesssim 1$ Gyr. This suggests a timeline for radio activation post detachment if the interaction zone lies at a polar latitude, with the reverse applying near the equator if magnetic flux is compressed there. Resistive relaxation simulations would help elucidate such interactions and their connections with diamagnetic screening, crystallisation, and convection.

### v4

@@ 3 Orbital gravitational waves
Figure 1: Characteristic strains, equation (8), for LPTs listed in Tab. 1 (see legends) assuming $P=P_{\text{orb}}$ and $T=2$ yr. Objects are color-coded according to different scenarios corresponding to chirp masses $\mathcal{M}$ set by expressions (3) (green), (4) (red), or (6) (blue). Exceptions are CHIME J0630 (Sec. 2.2.1), GPM J1839 (Sec. 2.2.2), and J1634 (Sec. 2.3.1), where orbital braking and Roche-lobe maxima restrict the component masses and companion makeup, respectively, if there is no active mass transfer. We consider two orbital periods for J1634, as detailed in Sec. 2.3.1. Filled symbols correspond to cases which exceed or overlap with the LISA or Taiji strain curves, taken from data provided by Robson, Cornish & Liu (2019) and Luo et al. (2020), respectively. The noise curve for Taiji is shown instead for the mission lifetime of 5 years to avoid visual clutter. Error bars on A J1448 are large due to the lack of distance constraints (Anumarlapudi et al., 2025).
@@
Assuming that the pulse period is that of the orbit, we use the data collated in Table 1 to estimate $h_{0}$ from equation (7) for each LPT using the chirp mass spread considered in Sec. 2. Notably, the period derivative is moderate in most systems ($|\dot{P}|_{\rm max}\lesssim 10^{-9}\text{ ss}^{-1}$; Coti Zelati & Borghese, 2024), meaning that GW signals should be essentially monochromatic over multi-year observational windows, $T$ (i.e., the folding time). This allows for an accumulation of signal power over many orbital cycles, $N=Tf_{\rm GW}=2T/P_{\rm orb}$. Averaging over orbital orientations and polarizations, the characteristic strain $h_{c}$ felt by a detector can thus be estimated through $h_{c}\approx\sqrt{2N}h_{0}$ (Finn & Thorne, 2000), with value
[…]
$\qquad\times\left(\frac{T}{0.5\text{ yr}}\right)^{1/2}\left(\frac{1\text{ kpc}}{d}\right).$
@@
Using the LISA design specifications from Amaro-Seoane et al. (2017), we adopt the noise power spectral density ($S_{n}$) fits from Robson, Cornish & Liu (2019), which include intrinsic noise (from single-link optical metrology and test mass acceleration) as well as Galactic confusion noise from unresolved sources, to estimate LPT detectability. For Taiji, the same Galactic confusion noise profile is used, but with the design specifications provided by Luo et al. (2020) and Chen & Liu (2025). Figure 1 shows characteristic strains $h_{c}$, calculated using equation (8), as functions of GW frequency for the objects listed in Tab. 1, assuming an observation time of two years. We find that three of the known LPTs could be detectable by LISA and Taiji (filled symbols), with several others being marginal under favorable conditions (see Sec. 3.1). It is noteworthy that the most promising candidates are also among the most intriguing.
@@
CHIME J0630. Given its proximity to Earth, insights on local supernovae rates could be gleaned if this source is a neutron star, a position which would be firm if no detection is made shortly after LISA launch. The period would thus correspond to rotation, requiring that some magnetars spindown faster than stipulated by pure dipole braking (e.g. via mass-loaded winds; Beniamini, Wadiasingh & Metzger, 2020). An excess of nearby neutron stars (and magnetars in particular) could also call for revisions to population synthesis models (though see Popov et al., 2003). While a nearby white dwarf is not particularly unusual, Tauris (2018) estimates that for S/Ns $\gtrsim 100$, $\mathcal{M}$ could be measured to within $\lesssim 1\%$ from which constraints on binary makeup, and possibly even pre-Newtonian effects predicted by some modified theories of gravity (Littenberg & Yunes, 2019), follow.
ILT/CHIME J1634. Given both its unusual polarization properties and confidently negative $\dot{P}$, this source is a compelling candidate for a binary. The orbital frequency is somewhat uncertain, and multiple scenarios with double-degenerates are viable (Dong et al., 2025b; Bloot et al., 2025). A joint LISA/Taiji detection would narrow down the possibilities by not only measuring the chirp mass but $f_{\rm GW}$ directly. Given that microstructure was observed within the pulses resembling that from XTE J1810–197 and the long-period pulsar PSR J0901–4046 (Caleb et al., 2022; Dong et al., 2025b), it is possible that this system consists of a tight binary with a magnetar. A detection may thus shed light on binary interactions in systems with strong magnetic fields (see, e.g., Glampedakis & Suvorov, 2021). The system should also be detectable $\text{S/N}(4\text{ yr)}\geq 10$] if $P_{\rm orb}$ corresponds instead to the longest ‘burst-like’ emissions ($P\sim 4036$ s) for chirp masses $\mathcal{M}\gtrsim 1M_{\odot}$.
@@
GX J1627. Discovered in 2018 archival data from the Murchison Widefield Array, this object has since remained silent (Hurley-Walker et al., 2022). If interpreted as a neutron star, because the spindown luminosity, $L_{\rm sd}\lesssim 10^{28}$ erg/s, is much lower than the radio luminosity, $L^{1.4\text{GHz}}_{\nu}\sim 4\times 10^{31}$ erg/s, it could not be rotation powered. If magnetically powered, the deep X-ray limits set by Chandra ($L^{\rm max}_{\rm X}\sim 10^{30}$ erg/s; Rea et al., 2022) imply that a non-detection by LISA could indicate that (some) LPTs populate a magnetar branch distinct from soft gamma repeaters and anomalous X-ray pulsars (Suvorov, Dehman & Pons, 2025). For a dwarf, the lack of radio reactivation suggests either a choked magnetosphere (e.g. due to accretion) or inconsistently beamed emissions, either of which could provide insight into tight-binary interactions.
@@
DA J1832. Because this LPT has shown contemporaneous X-ray bursting with radio pulsing (Wang et al., 2025), a magnetar progenitor is favored over a binary (Suvorov, Dehman & Pons, 2025). If confirmed as a binary however, sporadic accretion could be responsible for igniting a burst of comparable luminosity to that seen from DA J1832 in February 2024 ($L_{\rm X}\sim 10^{33}$ erg/s), though it would be difficult to explain the lack of any quiescent emissions (see Mukai, 2017). If the object enters into an outburst phase in future, searches for absorption lines could be useful to constrain its nature. The fact that the source was radio-loud for a $\sim$ year is also uncharacteristic for transitional pulsars (for instance), and so a detection could insist on revisions to theoretical models from a variety of directions.
@@
A J1935. This source displays a unique property amongst LPTs, being that of state switching: there are pulse states with high-degrees of linear polarization, others which are predominantly circular, and a null state (Caleb et al., 2024). A high degree of linear polarization and frequent nulling is rather reminiscent of radio-loud magnetars. For an isolated object, a highly-dynamic magnetosphere would be required to explain this temperamental behavior; the evolution could be driven, for example, by helicity injections from the plastic motions of crustal platelets (e.g. Beloborodov, 2009).
@@ 3.1 Signal-to-noise ratios
Table 2: S/Ns for LISA for the systems listed in Tab. 1 for $T=0.5$ yr (middle column) or $T=4$ yr (right) in either the least (largest distances and smallest chirp masses) or most (bracketed values) optimistic scenarios.
Source name S/N (0.5 yr; max) S/N (4 yr; max)
@@
In general, a confident detection requires an S/N of at least $\sim$ 10. While detection odds improve for sources with known sky locations, we provide conservative estimates by avoiding complications related to barycentric and other corrections. Borrowing formula (26) from Robson, Cornish & Liu (2019), the orientation- and polarization-averaged S/N is estimable through
$\text{S/N}\approx\frac{8\pi^{2/3}G^{5/3}\mathcal{M}^{5/3}f_{\rm GW}^{2/3}}{\sqrt{5}c^{4}d}\sqrt{\frac{T}{S_{n}(f_{\rm GW})}}\ .$ (9)
We calculate expression (9) for sources listed in Tab. 1. The results are shown in Table 2 for two scenarios: (i) using the maximum inferred distances (through dispersion measures) and lowest values of $\mathcal{M}$ considered in Sec. 2, and (ii) an optimistic case where the masses and distances take their upper and lower confidence bounds, respectively. The latter scenario is indicated by bracketed values in the second ($T=0.5$ yr) and third ($T=4$ yr) columns of Tab. 2.
As expected visually from Fig. 1, taking $T\lesssim 0.5$ yr practically guarantees that CHIME J0630 and GX J1627 would be identifiable in the LISA data stream if $P=P_{\rm orb}$. While hopes for other LPTs are low, there is some call for optimism for DA J1832 and A J1935 if the systems contain heavier stars, allowing them to reach significant S/Ns when accounting for their sky locations. For instance, if A J1935 contains a neutron star and a reasonably heavy companion, its well-constrained sky location could allow for a boost to the S/N by a factor $\sim 2$, which would give a final value of $\text{S/N}(4\text{yr})\sim 12$: sufficient to claim a confident detection.
3.2 Improvements via a network of detectors
Combining data from LISA with Taiji (and/or TianQin) could also lead to non-negligible increases in the S/N for each source, perhaps enabling detections in marginal scenarios like a WDWD for DA J1832. If we assume that the response of a network to some impinging GW is phase coherent in each detector, we can estimate the joint S/N through (Finn, 2001)
$\text{S/N}\approx\sqrt{(\text{S/N})_{\rm LISA}^{2}+(\text{S/N})_{\rm Taiji}^{2}+(\text{S/N})_{\rm TianQin}^{2}}.$ (10)
As discussed by Chen & Liu (2025), given the similarities between the Taiji and LISA noise curves (see Fig. 1) we expect the individual S/N values to be comparable for a given source. The S/N measurable by TianQin will be lower for an LPT. The total S/N (10), for any given LPT, may thus increase by a factor $\gtrsim\sqrt{2}$ relative to the results quoted in Tab. 2 (see also Sec. 4).
If the location of a given source is pinpointed in future, further enhancements may be expected since (correlated) confusion noises can be reduced. We anticipate a factor $\sim 2$ improvement in this case (Schutz, 2011; Robson, Cornish & Liu, 2019). Although a thorough statistical analysis lies beyond the scope of this article, such an increase could bolster the S/N past 10 in several marginal cases (e.g., a binary neutron-star scenario for A J1935 or a WDWD scenario for A J1448), thereby enabling a confident detection.
Aside from the obvious benefits anticipated from equation (10), there are indirect benefits that a network would offer. Given that LISA and Taiji will be positioned with large separations in space while in heliocentric orbits, combining their data streams could allow for a source to be accurately localized over the course of a long observational campaign (in terms of both distance and solid angle; Ruan et al., 2020). This could be especially useful to reduce error boxes to help in identifying optical and X-ray counterparts. A network of detectors can also veto glitch events (Schutz, 2011). While detector glitches are less important for long-lived signals, a range of scenarios could disable a detector for a period of time. The impact of offline time for a given detector would be mitigated with a network.
4 Outlook for future surveys
We have thus-far considered the detection prospects of known LPTs. Given the rapidity with which the field is developing, additional sources are likely to be discovered in future.
Figure 2: Sky-averaged signal-to-noise ratio (color scale), as a function of chirp mass and orbital period, for hypothetical LPTs jointly-observed by LISA and Taiji. An observation time of $T=4$ yr and a distance of $d=1$ kpc are assumed. Redder shades indicate higher S/N values.
As is clear from Tab. 1, there is no obvious clustering of periods in LPTs at present and the known range spans over an order of magnitude. In fact, a candidate LPT with a period of $P\sim 3.7$ days, containing a K7V dwarf, was recently discovered (J180526–292953; Frail et al., 2025). Nevertheless, we can investigate LPT detectability in general terms. Figure 2 shows sky-averaged S/Ns for a joint observation spanning 4 years for sources located at a distance of 1 kpc. In the bottom-right corner, corresponding to large $\mathcal{M}$ and short $P_{\rm orb}$, the highest S/N values are obtained ($\sim 10^{4}$). Such cases may apply to systems like GX J1627 in optimistic scenarios involving a neutron star (see Tab. 2), with leftward shifts for a WDWD or WDMD. The overlaid white curve marks $\text{S/N}=10$: LPTs below this line are likely detectable (Robson, Cornish & Liu, 2019). For example, for a WDWD with $\mathcal{M}=0.5M_{\odot}$ from (4), orbital periods of $P_{\rm orb}\lesssim 41$ min (similar to that of DA J1832) would be needed for LISA+Taiji to detect a signal. If the source were located instead at a J0630-like distance of $d=100$ pc, periods reaching $P_{\rm orb}\approx 114$ min would still enable $\text{S/N}>10$. If a $P\mathcal{M}$ correlation reveals itself in future, it may be possible to hunt for LPTs even without pulse data. In particular, candidates could be identified via GWs and localized with a network (Ruan et al., 2020), which could be used to guide radio surveys.
The leading explanation for pulsations in binary LPTs involves a relativistic cyclotron maser (Qu & Zhang, 2025). This mechanism requires the formation of a unipolar inductor, which will also work to drain orbital angular momentum together with GW losses (5) (e.g. Lai, 2012). The energy dissipation rate for GWs,
$\dot{E}_{\rm GW}\approx 10^{32}\left(\frac{\mathcal{M}}{0.4M_{\odot}}\right)^{10/3}\left(\frac{60\text{ min}}{P_{\rm orb}}\right)^{10/3}\text{erg/s},$ (11)
is typically large with respect to observed radio luminosities (e.g. $L_{\nu}\sim 10^{31}$ erg/s for the brightest single peaks in GX J1627; Hurley-Walker et al., 2022). Since perfectly-efficient conversion from electromagnetic dissipation to radio emission is unlikely though, some or even most of the orbital decay could be attributed to an inductor (Willes & Wu, 2004). A detection of $\dot{P}$, together with $\mathcal{M}$, could be used to assess this and implicitly constrain the radio emission mechanism. If the $\dot{P}$ and $\mathcal{M}$ values match closely with equation (5), pulsations would have to instead be fuelled by one of the individual objects. This could provide a way to identify magnetars in binaries, none of which are known to date (though see Suvorov & Glampedakis, 2022, for a discussion on LS I+61°303). By contrast, active mass transfer can widen the orbit and counterbalance GW losses. Future data could thus also be used to implicitly constrain the mass accretion rate to see if/how LPTs fit within the population of radio-loud catacyslmic variables (Ridder et al., 2023).
@@ 5 Conclusions
LPTs are mysterious, with some studies favoring isolated neutron stars and others favoring WDMDs or compact binaries (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). We have argued here that if the pulsation period in LPTs matches, or is close to, the orbital period, then at least a few should be visible to LISA and its sister spacecraft (Tab. 2). If electromagnetic data remain inconclusive, GW observations could help unveil their nature. Importantly, among the known LPTs, those that may be brightest in the LISA band correspond to objects with uncertain classification (Fig. 1).
Although we have presented a number of progenitor scenarios, the value of $\mathcal{M}$ is unknown in binary LPTs and must be searched over in a LISA/Taiji survey. This also applies to the distance (see Price, Flynn & Deller, 2021), meaning it is the combination $\mathcal{M}^{5/3}d^{-1}$ featuring in expression (9) that is particularly uncertain. Given the precision with which the period can/has been measured due to the timing of pulses though, it is unlikely that another source could be confused for a given LPT in the respective data streams upon Fourier binning (as the bin size $\Delta f\sim 1/T\ll f_{\rm GW}$). This applies especially to coherent searches, which are best-suited to reject coincident background signals (Klimenko et al., 2005). As highlighted at the beginning of Sec. 3, the period derivative is gradual in LPTs. This means that the signal should be essentially monochromatic and it is thus straightforward to construct many phase templates to carry out matched filtering (Schutz, 2011). In general, the uncertainty in $\dot{f}_{\rm GW}$ can be estimated as $\Delta\dot{f}_{\rm GW}\approx 4.3/(T^{2}\times\text{S/N})$ (Takahashi & Seto, 2002), from which the chirp mass uncertainty can be deduced as $\Delta\mathcal{M}/\mathcal{M}\approx 3/5\times\Delta\dot{f}_{\rm GW}/\dot{f}_{\rm GW}$ (Lau et al., 2020). We refer the interested reader to the above references for discussions on parameter surveys and search strategies.
While some implications of (non-)detections have been highlighted, a wealth of physical processes may operate in LPTs. For instance, linking radio luminosities to electromotive losses implies a minimum strength for the primary’s magnetic field. Assuming a maser-like emission mechanism, some systems appear to require GigaGauss strengths (Qu & Zhang, 2025). While dynamo action may amplify an internal field (Schreiber et al., 2021), the magnetospheric field may not reflect the interior strength because magnetic burial – a process where field lines are equatorially advected by infalling plasma – will reduce the global dipole moment during accretion while creating strong multipoles (Suvorov & Melatos, 2020). Typical accretion rates in tight-binary polars are $\sim 10^{-10}M_{\odot}\text{ yr}^{-1}$ (Pala et al., 2020) implying that, for a C/O dwarf with $M_{\star}\sim 0.8M_{\odot}$, the accretion timescale matches that of Ohmic diffusion ($\tau_{\Omega}$) at depths corresponding to densities of $\sim 10^{5}\text{ g cm}^{-3}$ where Cumming (2002) estimates $\tau_{\Omega}\lesssim 1$ Gyr. This suggests a timeline for radio activation post detachment if the interaction zone lies at a polar latitude, with the reverse applying near the equator if magnetic flux is compressed there. Resistive relaxation simulations would help elucidate such interactions and their connections with diamagnetic screening, crystallization, and convection.